# Characterizing Dynamical Stability of Stochastic Gradient Descent in Overparameterized Learning

### v2

@@ Abstract
For overparameterized optimization tasks, such as the ones found in modern machine learning, global minima are generally not unique. In order to understand generalization in these settings, it is vital to study to which minimum an optimization algorithm converges. The possibility of having minima that are unstable under the dynamics imposed by the optimization algorithm limits the potential minima that the algorithm can find. In this paper, we characterize the global minima that are dynamically stable/unstable for both deterministic and stochastic gradient descent (SGD). In particular, we introduce a characteristic Lyapunov exponent which depends on the local dynamics around a global minimum and rigorously prove that the sign of this Lyapunov exponent determines whether SGD can accumulate at the respective global minimum.
Keywords: machine learning, stochastic gradient descent, overparametrization, linear stability, Lyapunov exponents
@@ 1 Introduction
Since the success of “AlexNet” (Krizhevsky et al. (2012)), the overparameterization paradigm has become ubiquitous in modern machine learning. While it is well known that artificial neural networks with sufficiently many parameters can approximate arbitrary goal functions (see e.g. Hornik et al. (1989)), the training process, especially for overparameterized networks, is not well understood and leaves many open questions. For one, the fact that the loss landscapes are usually non-convex, makes it difficult to rigorously guarantee that the optimization algorithms converge to global minima.
@@
Another key open problem in understanding the training of overparameterized networks is that of generalization (see e.g. Zhang et al. (2021)). Classical wisdom suggests that overparameterized networks should suffer from overfitting problems due to their high expressivity. In practice, however, this is not the case. Put shortly, sufficiently large (i.e. deep, wide or both) artificial neural networks are capable of fitting arbitrary data, due to their large number of parameters (typically weights and biases). Thus, finding a set of parameters for which the network interpolates the training data does not in itself indicate that the network makes reasonable predictions for inputs outside the training data set. However, the parameters found by common optimization algorithms tend to perform well on unseen data. This is essential for the success of modern machine learning.
@@
One possible explanation for this phenomenon is that of dynamical stability (see e.g Wu et al. (2018)). Since the learning rate is in practice not infinitesimally small, some global minima can become dynamically unstable, in the sense that the optimization algorithm will not converge to these solutions, even if it is initialized arbitrarily close to them. This excludes unstable optimal solutions from being found by the optimization algorithm, thereby reducing the effective hypothesis class to stable global minima.
@@
For gradient descent, a second order Taylor expansions shows that a global minimum is only stable if the largest eigenvalue of the Hessian of the empirical loss is less than $2/\eta$, where $\eta$ is the learning rate. Thus gradient descent can only converge to minima which are sufficiently flat, which has been been associated to good generalization (see e.g. Hochreiter and Schmidhuber (1997)). Moreover, numerical experiments by Cohen et al. (2020) have demonstrated that (deterministic) gradient descent operates at the edge of stability, meaning that the largest eigenvalue of the Hessian oscillates around the critical value $2/\eta$ for the later stages of training.
@@
Transferring these ideas to the stochastic case requires an appropriate notion of dynamical stability for stochastic gradient descent (cf. Wu et al. (2018) and Ma and Ying (2021)). The goal of this work is to establish a coherent mathematical framework in which the dynamical stability of deterministic and stochastic optimization algorithms can be investigated and the convergence properties of global minima can be characterized.
@@ 1.1 Contributions
  • For global minima $x^{*}$, we introduce a new quantity, denoted by $\lambda(x^{*})$ (cf. (2.12) below), which can be seen as a characteristic Lyapunov exponent (Oseledec (1968)) and corresponds to a new notion of dynamic stability/instability for SGD. It should be noted that this notion differs [^1] In particular our notion of linar stability is strictly weaker than the one introduced in (Wu et al., 2018, Defintion 1) (cf. Appendix A). from the ones introduced in Wu et al. (2018) and Ma and Ying (2021) (cf. Appendix A).
@@
This clarifies the concept of dynamical stability for stochastic gradient descent and establishes a connection to a substantial body of mathematical and physical literature on Lyapunov exponents (see e.g. Oseledec (1968); Bougerol and Lacroix (1985); Arnold (1998)). Moreover, it provides a foundation for new theoretic and numerical investigations of the learning process under stochastic gradient descent. Possible questions of interest include how the value of $\lambda$ evolves during training with stochastic gradient descent and which implications a small, respectively large, value of $\lambda$ has on generalization.
From a mathematical perspective our analysis of SGD amounts to a study of the asymptotic behavior of discrete-time random dynamical systems (see e.g. Arnold (1998) and also section 3.5) possessing a manifold of equilibria. In the unstable case, we use a method based on moment Lyapunov exponents (see e.g. Arnold et al. (1986)). A more detailed discussion of this approach is given in section 3.1 an throughout the proofs.
@@ 1.2 Related Literature
Gurbuzbalaban et al. (2021) consider Lyapunov exponents of SGD in a setting where SGD does not converge to a global minimum, but the transition densities approach a heavy-tailed stationary distribution (see also e.g. Hodgkinson and Mahoney (2021)). Convergence rates for SGD with decaying learning rate have been studied in Fehrman et al. (2020).
The edge of stability phenomenon for gradient descent mentioned in the introduction has been numerically observed by Wu et al. (2018), as well as Cohen et al. (2020) and theoretically studied by Arora et al. (2022). Moreover, the difference between global minima found by GD and SGD (see Keskar et al. (2017); Wu et al. (2018)) and SGD and Adam (see Keskar and Socher (2017)) has been investigated numerically.
The dynamics of stochastic gradient descent has also been studied using approximations by stochastic differential equations, called stochastic modified equations (see e.g. Li et al. (2017, 2019)). We note that such an approach is not compatible with our analysis, as it assumes an infinitesimally small learning rate. Moreover, stochastic modified equations have been extended to a stochastic flow in Gess et al. (2024), building a framework analogous to the one we introduce in section 3.5.
The most critical part of our proof is based on techniques originally developed to study synchronization in stochastic differential equations by Baxendale and Stroock (1988) (see also Baxendale (1991)), which has seen recent attention motivated by fluid dynamics (Coti Zelati and Hairer (2021); Bedrossian et al. (2022); Blumenthal et al. (2023)).
@@ 1.3 Organization of the Paper
We start by introducing the fundamental setting of the optimization task (section 2.1), as well as the learning algorithms we consider (2.2). In section 2.3, we heuristically derive the notions of linear stability for GD and SGD. These derivations are made rigorous in our main results, which we present in section 2.4. In section 2.5, the main quantities of interest, $\mu$ and $\lambda$, are related to the neural tangent kernel (cf. Jacot et al. (2018)). Finally, we discuss possible generalizations of our main results in section 2.6. Section 3 provides the proofs of the main theorems. An overview over the structure of the proofs is given in section 3.1.
@@ 2.1 Network Model
In the following, we consider a scalar regression problem. Let $\hat{f}:\mathbb{R}^{d}\to\mathbb{R}$ be a ground-truth function, which is supposed to be reconstructed from $N$ given data pairs $(y_{i},z_{i}=\hat{f}(y_{i}))_{i\in[N]}$, where $[N]=\{1,\dots,N\}$. To do so, we consider a parameterized network model given by a smooth function [^2] For the purpose of being general, we will not specify how the network function $\mathfrak{F}$ looks like. Typical network architectures used in practice include fully connected networks, convolutional neural networks (image classification) and transformers (large language models). $\mathfrak{F}:\mathbb{R}^{D}\times\mathbb{R}^{d}\to\mathbb{R}$ and try, using the given data, to find an $x\in\mathbb{R}^{D}$ such that $\mathfrak{F}(x,\cdot)\approx\hat{f}$. Absent of other information, it is reasonable to prescribe
@@
for some loss-function :×$\ell:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{\geq 0}$, with $\ell(\bar{z},z)=0$ if and only if $\bar{z}=z$, $\partial_{1}\ell(z,z)=0$ and $\partial_{1}^{2}\ell(z,z)=1$. Even though other choices for $\ell$ are possible, it is advised to think of the square loss function $\ell(\bar{z},z):=\frac{1}{2}(\bar{z}-z)^{2}$. For each $i\in[N]$, we define the individual loss function $\mathcal{L}_{i}:\mathbb{R}^{D}\to[0,\infty)$ by
@@
and the empirical loss function $\mathcal{L}:\mathbb{R}^{D}\to\mathbb{R}$ as the average of the individual loss functions, i.e.
@@
If the number of given training examples $N$ exceeds the number of learnable parameters $D$, the problem is overdetermined and the global minimum of $\mathcal{L}$ is usually unique, such that $x$ is fully determined by (2.4). However, in general it will not be possible to find an $x\in\mathbb{R}^{D}$ such that $\mathfrak{F}(x,\cdot)$ interpolates the given data, i.e. the set
@@
As a consequence of Sard’s theorem (see e.g. Milnor (1997)), this condition will be satisfied for generic training data [^3] Put more precisely, for every smooth network model $\mathfrak{F}$ and every choice of $(y_{1},\dots,y_{N})$, there exists a set $\mathcal{Z}\subseteq\mathbb{R}^{N}$ of full Lebesgue measure, such that, if $(z_{1},\dots,z_{N})\in\mathcal{Z}$, then $\mathcal{M}$ is an embedded smooth $(D-N)$-dimensional manifold (cf. Cooper (2021)). and in the following we will assume that $\mathcal{M}$ is a smooth manifold as described (cf. Hypothesis (H1) below).
@@ 2.2 Learning Algorithms
In practice, a wide range of optimization algorithms (cf. e.g. Schmidt et al. (2021)) are used to find a global minimum $x^{*}\in\mathcal{M}$. In this work, we will restrict our study to the two most basic optimization algorithms: gradient descent (GD) and stochastic gradient descent (SGD). Both algorithms are initialized at a random point $X_{0}^{\operatorname{GD/SGD}}\in\mathbb{R}^{D}$, which is distributed according to some probability measure $\nu:\mathfrak{B}(\mathbb{R}^{D})\to[0,1]$, called initial distribution. [^4] Here and in the following, $\mathfrak{B}(\mathbb{R}^{D})$ denotes the family of Borel sets. We will assume that $\nu$ is equivalent to Lebesgue measure [^5] This is for example consistent with $\nu$ being a normal distribution, which is the most common choice in practice. (cf. Hypothesis (H2) below).
Gradient descent tries to successively reduce the empirical loss $\mathcal{L}$ by taking small steps in the opposite direction of its gradient $\nabla\mathcal{L}$. Formally the update rule is given by
@@
where $\eta>0$ is a small real number called learning rate. Ideally, the hope is that gradient descent converges to some point ${X^{\operatorname{GD}}_{\lim}\in\mathcal{M}}$ as $n$ tends to infinity. While gradient descent is remarkably reliable at finding global minima in practice, convergence to a global minimum (or even convergence in the first place) can not be guaranteed in our general setting, since $\mathcal{L}$ is generally non-convex. Therefore we define $X^{\operatorname{GD}}_{\lim}$ as an ${}\mathcal{M}\cup\{\emptyset\}$-valued random variable by
@@
accounting for a possible failure of convergence. Note that the randomness in $X^{\operatorname{GD}}_{\lim}$ stems purely from the random initialization $X^{\operatorname{GD}}_{0}$. Once $X^{\operatorname{GD}}_{0}$ is drawn all subsequent steps and thus $X^{\operatorname{GD}}_{\lim}$ are deterministic.
@@ 2.3 Linear Stability
and noting that the right-hand sides vanish if $x\in\mathcal{M}$. Since the initial distribution $\nu$ is assumed to have full support, it is a priori possible that $X^{\operatorname{GD/SGD}}_{\lim}$ takes any value in $\mathcal{M}$. However, under mild assumptions (cf. (H2) and (H3) below), the optimization algorithm can only reach a global minimum in a finite number of steps with probability zero. In other words, (S)GD will usually converge to some global minimum, say $x^{*}\in\mathcal{M}$, without reaching it in a finite number of steps. This is only possible, if $x^{*}$ is dynamically stable, meaning that, once the optimization algorithms gets sufficiently close to $x^{*}$, it will stay close.
In order to study dynamical stability at some $x^{*}\in\mathcal{M}$, we linearize the optimization step (2.7) around $x^{*}$. We introduce the function $\mathcal{G}_{\eta}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ by
@@
Note that, by definition of the normal space $\mathcal{N}$ (2.5) and tangent space $\mathcal{T}$ (2.6), the matrix $\mathcal{G}^{\prime}_{\eta}(x^{*})$ satisfies $\mathcal{G}^{\prime}_{\eta}(x^{*})v=v\in\mathcal{T}(x^{*})$ for all $v\in\mathcal{T}(x^{*})$ and $\mathcal{G}^{\prime}_{\eta}(x^{*})w\in\mathcal{N}(x^{*})$ for all $w\in\mathcal{N}(x^{*})$. Thus, $\mathcal{G}^{\prime}_{\eta}(x^{*})$ respects the splitting $\mathbb{R}^{D}=\mathcal{T}(x^{*})\oplus\mathcal{N}(x^{*})$ and the restriction $\mathcal{G}^{\prime}_{\eta}(x^{*})|_{\mathcal{T}(x^{*})}:\mathcal{T}(x^{*})\to\mathcal{T}(x^{*})$ equals the identity. We say that $x^{*}$ is linearly stable under GD with learning rate $\eta$ if
@@
where the last equivalence holds if $\eta<2$. While this formulation is more common in the literature (cf. e.g. Wu et al. (2018); Cohen et al. (2020); Arora et al. (2022)), we will stick to the expression $\mu(x^{*})<0$ here, as it can be more easily extended to stochastic gradient descent.
Our stability analysis for stochastic gradient descent is analogous to that for gradient descent. For notational convenience we introduce the functions $\mathcal{G}_{\eta,i}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ for each $i\in[N]$ by
@@
Note that the matrices $\mathcal{G}^{\prime}_{\eta,i}(x^{*})$ also all respect the splitting $\mathbb{R}^{D}=\mathcal{T}(x^{*})\oplus\mathcal{N}(x^{*})$ with $\mathcal{G}^{\prime}_{\eta,i}(x^{*})|_{\mathcal{T}(x^{*})}=\operatorname{Id}_{\mathcal{T}(x^{*})}$. Analogously to gradient descent, we would like to call $x^{*}$ linearly stable, if
@@
At first glance, this definition raises two problems. Firstly, it is not clear whether the limit exists. Secondly, the value of $\lambda(x^{*})$ seems to depend on the random sequence $(\xi_{1},\xi_{2},\dots)$. However, Kingman’s sub-additive ergodic theorem (Kingman (1968), see also Steele (1989)) [^6] The integrability condition for Kingman’s ergodic theorem holds trivially here, as we only consider finitely many different matrices. states that for almost every realization $(\xi_{1},\xi_{2},\dots)$ we have
@@
In fact, the following theorem due to Oseledec (1968) (see also e.g. chapter 3.4 in Arnold (1998)) shows the existence of so-called Oseledec’s subspaces which are the equivalent notion to generalized eigenspaces for random matrix products.
@@
In particular for every $y\in\mathcal{N}\setminus V_{2}(\xi_{1},\xi_{2},\dots)$, we have

### v3

@@ Abstract
For overparameterized optimization tasks, such as those found in modern machine learning, global minima are generally not unique. In order to understand generalization in these settings, it is vital to study to which minimum an optimization algorithm converges. The possibility of having minima that are unstable under the dynamics imposed by the optimization algorithm limits the potential minima that the algorithm can find. In this paper, we characterize the global minima that are dynamically stable/unstable for both deterministic and stochastic gradient descent (SGD). In particular, we introduce a characteristic Lyapunov exponent that depends on the local dynamics around a global minimum and rigorously prove that the sign of this Lyapunov exponent determines whether SGD can accumulate at the respective global minimum.
Keywords: machine learning, stochastic gradient descent, overparameterization, linear stability, Lyapunov exponents
@@ 1 Introduction
Since the success of “AlexNet” (Krizhevsky et al., 2012), the overparameterization paradigm has become ubiquitous in modern machine learning. Although it is well known that artificial neural networks with sufficiently many parameters can approximate arbitrary goal functions (see, e.g., Hornik et al., 1989), the training process, especially for overparameterized networks, is not well understood and leaves many open questions. For one, the fact that the loss landscapes are usually non-convex makes it difficult to rigorously guarantee that the optimization algorithms converge to global minima.
@@
Another key open problem in understanding the training of overparameterized networks is that of generalization (see, e.g., Zhang et al., 2021). Classical wisdom suggests that overparameterized networks should suffer from overfitting problems due to their high expressivity. However, in practice, this is not the case. In other words, sufficiently large (i.e. deep, wide, or both) artificial neural networks are capable of fitting arbitrary data, due to their large number of parameters (typically weights and biases). Thus, finding a set of parameters for which the network interpolates the training data, in itself, does not indicate that the network makes reasonable predictions for inputs outside the training data set. However, the parameters found by common optimization algorithms tend to perform well on unseen data. This is essential for the success of modern machine learning.
@@
One possible explanation for this phenomenon is that of dynamical stability (see, e.g. Wu et al., 2018). Since the learning rate is in practice not infinitesimally small, some global minima can become dynamically unstable, in the sense that the optimization algorithm will not converge to these solutions, even if it is initialized arbitrarily close to them. This excludes unstable optimal solutions from being found by the optimization algorithm, thereby reducing the effective hypothesis class to stable global minima.
@@
For gradient descent, a second-order Taylor expansion shows that a global minimum is only stable if the largest eigenvalue of the Hessian of the empirical loss is less than $2/\eta$, where $\eta$ is the learning rate. Thus, gradient descent can only converge to minima that are sufficiently flat, which has been associated with good generalization (see e.g. Hochreiter and Schmidhuber, 1997). Moreover, numerical experiments by Cohen et al. (2020) have demonstrated that (deterministic) gradient descent operates at the edge of stability, which means that the largest eigenvalue of the Hessian oscillates around the critical value $2/\eta$ for the later stages of training.
@@
Transferring these ideas to the stochastic case requires an appropriate notion of dynamical stability under stochastic gradient descent (cf. Wu et al., 2018 and Ma and Ying, 2021). The goal of this work is to establish a coherent mathematical framework in which the dynamical stability of deterministic and stochastic optimization algorithms can be investigated and the convergence properties of global minima can be characterized.
@@ 1.1 Contributions
  • For global minima $x^{*}$, we introduce a new quantity, denoted by $\lambda(x^{*})$ (cf. (2.12) below), which can be seen as a characteristic Lyapunov exponent (Oseledec, 1968) and corresponds to a new notion of dynamic stability/instability for SGD. It should be noted that this notion differs [^1] In particular our notion of linear stability is strictly weaker than the one introduced in (Wu et al., 2018, Defintion 1) (cf. Appendix A). from the ones introduced in Wu et al. (2018) and Ma and Ying (2021) (cf. Appendix A).
@@
This clarifies the concept of dynamical stability for stochastic gradient descent and establishes a connection to a substantial body of mathematical and physical literature on Lyapunov exponents (see, e.g. Oseledec, 1968; Bougerol and Lacroix, 1985; Arnold, 1998). Moreover, it provides a foundation for new theoretic and numerical investigations of the learning process under stochastic gradient descent. Possible questions of interest include how the value of $\lambda$ evolves during training with stochastic gradient descent and what implications a small, respectively large, value of $\lambda$ has on generalization.
From a mathematical perspective, our analysis of SGD amounts to a study of the asymptotic behavior of discrete-time random dynamical systems (see, e.g. Arnold, 1998 and also Section 3.5) possessing a manifold of equilibria. In the unstable case, we use a method based on moment Lyapunov exponents (see, e.g. Arnold et al., 1986). A more detailed discussion of this approach is given in Section 3.1 and throughout the proofs.
@@ 1.2 Related Literature
Gurbuzbalaban et al. (2021) consider Lyapunov exponents of SGD in a setting where SGD does not converge to a global minimum, but the transition densities approach a heavy-tailed stationary distribution (see also e.g. Hodgkinson and Mahoney, 2021). The convergence rates for SGD with decaying learning rate have been studied in Fehrman et al. (2020).
The edge of stability phenomenon for gradient descent mentioned in the introduction has been numerically observed by Wu et al. (2018), as well as by Cohen et al. (2020) and theoretically studied by Arora et al. (2022). Moreover, the difference between the global minima found by GD and SGD (see Keskar et al., 2017; Wu et al., 2018) and between SGD and Adam (see Keskar and Socher, 2017) has been numerically investigated.
The dynamics of stochastic gradient descent have also been studied using approximations by stochastic differential equations, called stochastic modified equations (see, e.g. Li et al. (2017, 2019)). We note that such an approach is not compatible with our analysis as it assumes an infinitesimally small learning rate. Moreover, the modified stochastic equations have been extended to a stochastic flow in Gess et al. (2024), building a framework analogous to the one we introduce in Section 3.5.
The most critical part of our proof is based on techniques originally developed to study synchronization in stochastic differential equations by Baxendale and Stroock (1988) (see also Baxendale, 1991), which has received recent attention motivated by fluid dynamics (Coti Zelati and Hairer, 2021; Bedrossian et al., 2022, 2025; Blumenthal et al., 2023).
@@ 1.3 Organization of the Paper
We start by introducing the fundamental setting of the optimization task (Section 2.1), as well as the learning algorithms that we consider (2.2). In Section 2.3, we heuristically derive the notions of linear stability for GD and SGD. These derivations are made rigorous in our main results, which we present in Section 2.4. In Section 2.5, the main quantities of interest, $\mu$ and $\lambda$, are related to the neural tangent kernel (cf. Jacot et al., 2018). Finally, we discuss possible generalizations of our main results, such as extensions to other optimizers, in Section 2.7. Section 3 provides the proofs of the main theorems. An overview of the structure of the proofs is given in Section 3.1.
@@ 2.1 Network Model
In the following, we consider a scalar regression problem. Let $\hat{f}:\mathbb{R}^{d}\to\mathbb{R}$ be a ground truth function, which is supposed to be reconstructed from $N$ given data pairs $(y_{i},z_{i}=\hat{f}(y_{i}))_{i\in[N]}$, where $[N]=\{1,\dots,N\}$. To do so, we consider a parameterized network model given by a smooth function [^2] For the purpose of being general, we will not specify how the network function $\mathfrak{F}$ looks like. Typical network architectures used in practice include fully connected networks, convolutional neural networks (image classification), and transformers (large language models). $\mathfrak{F}:\mathbb{R}^{D}\times\mathbb{R}^{d}\to\mathbb{R}$ and try, using the given data, to find an $x\in\mathbb{R}^{D}$ such that $\mathfrak{F}(x,\cdot)\approx\hat{f}$. In the absence of other information, it is reasonable to prescribe
@@
for some loss function :×$\ell:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{\geq 0}$, with $\ell(\bar{z},z)=0$ if and only if $\bar{z}=z$, $\partial_{1}\ell(z,z)=0$ and $\partial_{1}^{2}\ell(z,z)=1$. Although other choices for $\ell$ are possible, it is recommended to think of the square loss function $\ell(\bar{z},z):=\frac{1}{2}(\bar{z}-z)^{2}$. For each $i\in[N]$, we define the individual loss function $\mathcal{L}_{i}:\mathbb{R}^{D}\to[0,\infty)$ by
@@
and the empirical loss function $\mathcal{L}:\mathbb{R}^{D}\to\mathbb{R}$ as the average of the individual loss functions, that is,
@@
If the number of given training examples $N$ exceeds the number of learnable parameters $D$, the problem is overdetermined and the global minimum of $\mathcal{L}$ is usually unique, so that $x$ is fully determined by (2.4). However, in general, it will not be possible to find an $x\in\mathbb{R}^{D}$ such that $\mathfrak{F}(x,\cdot)$ interpolates the given data, i.e. the set
@@
As a consequence of Sard’s theorem (see e.g. Milnor, 1997), this condition will be satisfied for generic training data [^3] Put more precisely, for every smooth network model $\mathfrak{F}$ and every choice of $(y_{1},\dots,y_{N})$, there exists a set $\mathcal{Z}\subseteq\mathbb{R}^{N}$ of full Lebesgue measure, such that, if $(z_{1},\dots,z_{N})\in\mathcal{Z}$, then $\mathcal{M}$ is an embedded smooth $(D-N)$-dimensional manifold (cf. Cooper, 2021). and in the following we will assume that $\mathcal{M}$ is a smooth manifold as described (cf. Hypothesis (H1) below).
@@ 2.2 Learning Algorithms
In practice, a wide range of optimization algorithms (cf. e.g. Schmidt et al., 2021) are used to find a global minimum $x^{*}\in\mathcal{M}$. In this work, we will restrict our study to the two most basic optimization algorithms: gradient descent (GD) and stochastic gradient descent (SGD). Both algorithms are initialized at a random point $X_{0}^{\operatorname{GD/SGD}}\in\mathbb{R}^{D}$, which is distributed according to some probability measure $\nu:\mathfrak{B}(\mathbb{R}^{D})\to[0,1]$, called the initial distribution. [^4] Here and in the following, $\mathfrak{B}(\mathbb{R}^{D})$ denotes the family of Borel sets. We will assume that $\nu$ is equivalent to Lebesgue measure [^5] This is for example consistent with $\nu$ being a normal distribution, which is the most common choice in practice. (cf. Hypothesis (H2) below).
Gradient descent tries to successively reduce the empirical loss $\mathcal{L}$ by taking small steps in the opposite direction of its gradient $\nabla\mathcal{L}$. Formally, the update rule is given by
@@
where $\eta>0$ is a small real number called learning rate. Ideally, the hope is that the gradient descent converges to some point ${X^{\operatorname{GD}}_{\lim}\in\mathcal{M}}$ as $n$ tends to infinity. Although gradient descent is remarkably reliable in finding global minima in practice, convergence to a global minimum (or even convergence in the first place) cannot be guaranteed in our general setting, since $\mathcal{L}$ is generally non-convex. Therefore we define $X^{\operatorname{GD}}_{\lim}$ as an ${}\mathcal{M}\cup\{\emptyset\}$-valued random variable by
@@
accounting for a possible failure of convergence. Note that the randomness in $X^{\operatorname{GD}}_{\lim}$ stems purely from the random initialization $X^{\operatorname{GD}}_{0}$. Once $X^{\operatorname{GD}}_{0}$ is drawn, all subsequent steps, and thus $X^{\operatorname{GD}}_{\lim}$ are deterministic.
@@
$\operatorname{supp}\left(X^{\operatorname{GD/SGD}}_{\lim}\right):=\left\{x\in\mathcal{M}:{\mathbb{P}\left(X^{\operatorname{GD/SGD}}_{\lim}\in U\right)>0,~{}\forall\,U\subseteq\mathcal{M}\text{ open nbhd. of }x}\right\}.$
@@ 2.3 Linear Stability
and noting that the right-hand sides vanish if $x\in\mathcal{M}$. Since the initial distribution $\nu$ is assumed to have full support, it is possible that $X^{\operatorname{GD/SGD}}_{\lim}$ takes any value in $\mathcal{M}$. However, under mild assumptions (cf. (H2) and (H3) below), the optimization algorithm can only reach a global minimum in a finite number of steps with probability zero. In other words, (S)GD will usually converge to some global minimum, say $x^{*}\in\mathcal{M}$, without reaching it in a finite number of steps. This is only possible if $x^{*}$ is dynamically stable, which means that, once the optimization algorithm gets sufficiently close to $x^{*}$, it will stay close.
In order to study the dynamical stability at some $x^{*}\in\mathcal{M}$, we linearize the optimization step (2.7) around $x^{*}$. We introduce the function $\mathcal{G}_{\eta}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ by
@@
Note that, by definition of the normal space $\mathcal{N}$ (2.5) and the tangent space $\mathcal{T}$ (2.6), the matrix $\mathcal{G}^{\prime}_{\eta}(x^{*})$ satisfies $\mathcal{G}^{\prime}_{\eta}(x^{*})v=v\in\mathcal{T}(x^{*})$ for all $v\in\mathcal{T}(x^{*})$ and $\mathcal{G}^{\prime}_{\eta}(x^{*})w\in\mathcal{N}(x^{*})$ for all $w\in\mathcal{N}(x^{*})$. Thus, $\mathcal{G}^{\prime}_{\eta}(x^{*})$ respects the splitting $\mathbb{R}^{D}=\mathcal{T}(x^{*})\oplus\mathcal{N}(x^{*})$ and the restriction $\mathcal{G}^{\prime}_{\eta}(x^{*})|_{\mathcal{T}(x^{*})}:\mathcal{T}(x^{*})\to\mathcal{T}(x^{*})$ equals the identity. We say that $x^{*}$ is linearly stable under GD with learning rate $\eta$ if
@@
where the last equivalence holds if $\eta<2$. Although this formulation is more common in the literature (cf., e.g. Wu et al., 2018; Cohen et al., 2020; Arora et al., 2022), we will stick to the expression $\mu(x^{*})<0$ here, as it can be more easily extended to stochastic gradient descent.
Our stability analysis for stochastic gradient descent is analogous to that for gradient descent. For notational convenience, we introduce the functions $\mathcal{G}_{\eta,i}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ for each $i\in[N]$ by
@@
Note that the matrices $\mathcal{G}^{\prime}_{\eta,i}(x^{*})$ also all respect the splitting $\mathbb{R}^{D}=\mathcal{T}(x^{*})\oplus\mathcal{N}(x^{*})$ with $\mathcal{G}^{\prime}_{\eta,i}(x^{*})|_{\mathcal{T}(x^{*})}=\operatorname{Id}_{\mathcal{T}(x^{*})}$. Analogously to gradient descent, we would like to call $x^{*}$ linearly stable if
@@
At first glance, this definition raises two problems. First, it is not clear whether the limit exists. Secondly, the value of $\lambda(x^{*})$ seems to depend on the random sequence $(\xi_{1},\xi_{2},\dots)$. However, Kingman’s sub-additive ergodic theorem (Kingman, 1968, see also Steele, 1989) [^6] The integrability condition for Kingman’s ergodic theorem holds trivially here, as we only consider finitely many different matrices. states that for almost every realization $(\xi_{1},\xi_{2},\dots)$ we have
@@
In fact, the following theorem due to Oseledec (1968) (see also e.g. chapter 3.4 in Arnold, 1998) shows the existence of so-called Oseledec’s subspaces, which are the equivalent notion to generalized eigenspaces for random-matrix products.
@@
In particular, for every $y\in\mathcal{N}\setminus V_{2}(\xi_{1},\xi_{2},\dots)$, we have